Let f(x) = 3x^4 – 1. Which of the following is true?
A. f(-x)=-f(x)
B. f(-x)= f (x)
C. f(-x)=1/f(x)
D. None of the above

Answer by nerdybill(7384)

Let f(x) = 3x^4 – 1. Which of the following is true?
A. f(-x)=-f(x)
B. f(-x)= f (x)
C. f(-x)=1/f(x)
D. None of the above
.
set x to -x and evaluate:
f(x) = 3x^4 – 1
f(-x) = 3(-x)^4 – 1
f(-x) = 3(-1)^4(x)^4 – 1
f(-x) = 3(x)^4 – 1
f(-x) = 3x^4 – 1
So, answer is D -- None of the above